# Odo[Uncaptioned image]: Depth-Guided Diffusion for Identity-Preserving Body Reshaping

### v2

@@ 1 Introduction
Previous methods for human shape editing have primarily relied on 3D morphable models or image warping techniques. However, these methods often introduce visual artifacts and background distortions, especially under large deformations, due to their reliance on explicit geometric correspondences. In contrast, human pose editing has seen substantial progress, largely due to the availability of large-scale pose editing datasets [20]. Capturing diverse poses is straightforward, as an individual can rapidly shift poses during short capture sessions, or poses can be automatically extracted from video footage. However, gathering data of the same person exhibiting multiple distinct body shapes presents significant practical challenges. Thus, the absence of large-scale datasets with identity-consistent body shape variations remains a critical barrier to developing robust shape editing techniques.
To address this, we introduce ChangeLing18K, the first large-scale dataset specifically curated for human shape editing. It features diverse body types including thin, muscular, overweight, with each pair depicting the same subject across different shapes while maintaining consistent identity, pose, clothing, and viewpoint. This facilitates precise and identity-preserving model training. Building on this dataset, we propose Odo, to the best of our knowledge, the first ever end-to-end diffusion-based framework designed specifically for human shape editing. Recently, diffusion models have demonstrated remarkable success in different image-editing tasks [23, 6, 22, 16]. Their iterative denoising mechanism, combined with strong priors learned from large datasets, enables fine-grained control and high-quality outputs. However, traditional methods like text-based or point-dragging-based methods [3, 39] and are not capable of human body shape transformation. These techniques lack the precise control for editing human body shapes often causing artifacts or undesired results.
@@
Figure 2: Interactive UI with semantic sliders for shape editing. In this example, the semantic sliders are adjusted to modify the body shape, resulting in the person appearing significantly fatter compared to the original input image.
@@ 2 Related Work
Human body shape editing requires specialized approaches beyond general image editing, as it demands precise anatomical transformations of the human body. In this section we discuss the evolution of human-shape editing methods, highlight general diffusion-based techniques, and the current scope of SMPL conditioning in diffusion models.
@@
Image Editing using Diffusion. Foundational work such as Denoising Diffusion Probabilistic Models (DDPM) [11] and Latent Diffusion Models (LDM) [35] introduced scalable frameworks that enable controllable and high-resolution image synthesis. These models have served as the foundation for numerous diffusion-based image-editing works across a broad range of tasks, including inpainting, colorization, semantic modification, and subject personalization.
@@
These methods are designed for semantic, structural, or stylistic editing but are unsuitable for large-scale anatomical transformations in the human body while preserving fine-grained identity and clothing details. Human shape editing requires manipulating body proportions in a controllable manner, beyond the capacity of latent inversion or prompt-conditioned editing.
@@
SMPL Conditioned Diffusion. Recent methods integrate SMPL models into diffusion frameworks to enhance spatial and anatomical consistency in human image generation. CHAMP [48] utilizes SMPL-derived spatial representations as conditioning inputs to a latent diffusion model. By aligning these SMPL conditions with a reference image’s shape parameters, CHAMP achieves temporally coherent human animations. However, its scope is limited to pose variation and does not support body shape editing. Buchheim et al. [4] propose an SMPL-conditioned ControlNet for text-to-image generation, explicitly encoding SMPL shape and pose parameters into the diffusion UNet cross-attention. However, their method requires synthetic-to-real domain adaptation and does not directly facilitate shape editing. These methods underscore the potential of SMPL conditioning to enhance human shape control within diffusion models.
3 ChangeLing18K Dataset
@@ 3.1 Dataset Generation Workflow
ChangeLing18K is synthesized through a four–stage, model–driven pipeline that guarantees identity fidelity, pose and attire consistency, and high–resolution realism across three body–types (thin, fat, muscular) Fig. 4. The complete dataset generation pipeline is shown in Fig. 3 and detailed below.
@@
Thin Fat Muscular
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: The left figure illustrates our training pipeline, where the SMPL depth-map extracted from the target image conditions the ReshapeNet via ControlNet for the desired shape transformation. Features from the reference image, extracted by the ReferenceNet, are integrated into the ReshapeNet using spatial self-attention. The right figure depicts the inference pipeline, where SMPL parameters are initialized from the input image, adjusted based on semantic attributes, and rendered into a target SMPL depth map. This depth map, along with input image features, conditions the ReshapeNet during inference to generate the final transformed output through iterative denoising.
@@
i) Reference Image selection. The thin image is treated as the baseline reference because it’s smaller silhoutte exposes the most background, yielding a neear complete scene for subsequent background completion step.
@@
iv) Uniform Scaling & Composition. From the thin reference mask we measure the height and record the bottommost coordinates of the mask. Each non-thin mask is resized uniformly to match it’s height, maintaining aspect ratio, and composited onto the clean background at the same coordinates. This ensures consistent framing across images while preserving real shape variations.
@@ 4 Method
Given the above dataset we propose Odo, a diffusion-based approach for reshaping humans in images using target SMPL depth maps. The SMPL model [21] represents the human body’s 3D shape with parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ and pose with joint rotation parameters $\boldsymbol{\theta}\in\mathbb{R}^{72}$. To generate the conditioning depth maps, we first fit the SMPL model to each individual in our dataset using the method introduced in [37]. The optimized $\boldsymbol{\beta}$ and $\boldsymbol{\theta}$ parameters are subsequently used to render spatially aligned depth maps.
During training, we use paired images from our dataset of the same individual, captured with consistent clothing and background, but exhibiting different body shapes. Given an input image $\boldsymbol{I}$ and an SMPL target depth map $\boldsymbol{S}$, our objective is to transform the person in $\boldsymbol{I}$ to match the shape in $\boldsymbol{S}$.
For inference on in-the-wild images without ground-truth SMPL depth maps, we use a mapping between semantic attributes and SMPL $\boldsymbol{\beta}$ shape parameters [8]. This mapping enables modification of semantic attributes like weight, height, and body size to control the SMPL $\boldsymbol{\beta}$ shape parameters. Adjusting these parameters modifies the depth map, ensuring the generated output image aligns with the desired body shape. We show this interface in Fig. 2.
@@ 5.1 Implementation Details
We train Odo on our training set combined with a subset of the DeepFashion-MultiModal dataset [14], totaling 20,000 image pairs. The DeepFashion-MultiModal images provide exposure to complex garments and varied faces, with images paired with themselves during training. All SMPL depth maps are precomputed to minimize memory usage and enhance training efficiency.
@@
For initialization, the ReshapeNet uses pre-trained SDXL UNet weights and is fine-tuned jointly with the IP-Adapter (initialized from its pre-trained checkpoint). The ReferenceNet is initialized with SDXL weights and kept frozen during training. The ControlNet uses pre-trained depth ControlNet weights [42] and also remains frozen throughout training.
Our model requires approximately 23GB GPU memory and 18 seconds for single-image inference.
@@ 5.2 Evaluation
As noted by Okuyama et al. [28], no dataset exists for objective quantitative evaluation of body shape editing with large deformations. We therefore construct the first benchmark dataset for evaluating human body shape editing, consisting of 3,600 image pairs. The dataset uses distinct real face images and background descriptions to prevent training data leakage and includes diverse body shape variations for comprehensive assessment.
We evaluate generated images against ground truth using four key metrics. For image fidelity, we use Structural Similarity Index Measure (SSIM[41], Peak Signal-to-Noise Ratio (PSNR) for pixel-level accuracy, and Learned Perceptual Image Patch Similarity (LPIPS[46] for perceptual similarity. For body shape accuracy, we fit SMPL models to both ground truth and generated images using [37], then compute the Scale Corrected Per-Vertex Euclidean error in neutral (T-)pose (PVE-T-SC[38]. This metric quantifies how accurately the generated image preserves the target body shape.
@@
In-the-wild Images. Fig. 6 demonstrates inference on in-the-wild images, where our method effectively handles diverse poses, backgrounds, and clothing while preserving person identity. In addition to SMPL target shapes, we provide textual prompts—“Make the person fatter,” “Make the person thinner,” or “Make the person muscular”—to explicitly guide the desired transformations.
Fig. 1 further demonstrates our model’s capability to perform diverse shape transformations. The model accurately follows SMPL depth maps to generate multiple variations of thinner and fatter versions from the reference image.
@@
We also evaluate FLUX.1 Kontext[dev] [2], a prompt-based editing model. We design specific prompts instructing the model to “Make the person fatter,” “Make the person thinner,” or “Make the person muscular” while specifying target weights. However, due to the lack of fine-grained editing controls, FLUX.1 Kontext[dev] achieves limited performance, as shown in Tab. 2.
Table 2: Comparison of our model Odo against Ren et al. [33] and FLUX.1 Kontext[dev] on our test set. We also show results for three ablations: without prompt conditioning in ReshapeNet, without ReferenceNet (IP-Adapter only), and training only on BR-5K data.
@@ 5.3 Ablations
We conduct ablation experiments to evaluate the effectiveness of various modules in our model. Specifically, we investigate different methods for inputting image features into ReshapeNet and also explore different prompting techniques to guide the model in performing specific shape transformation tasks.
@@
We also train our model only on the BR-5K dataset and evaluate it on our benchmark data to highlight the significance of our new dataset. Since the BR-5K dataset includes only very slight variations in body shape, the model is unable to learn to perform any significant body shape transformations as seen in Fig. 8. This is also reflected in the high value of the per-vertex error metric obtained on the test set as shown in Tab. 2.
The quantitative results for these ablations are shown in Tab. 2.
@@ 6 Conclusion
We introduce Odo, the first end-to-end diffusion-based framework for photorealistic, identity-preserving body shape editing at high resolution. Our approach generalizes across diverse images with varying poses and backgrounds by combining a ReferenceNet with a Depth ControlNet-conditioned diffusion model to produce fine-grained and diverse body shape modifications. We also introduce ChangeLing18K, a large-scale human body reshape dataset containing 18,573 paired images with varied transformations while maintaining consistent identity, clothing, and backgrounds.
Despite these advances, our model has certain limitations. It occasionally struggles to preserve fine facial details, leading to identity loss and facial artifacts. Additionally, inaccuracies during SMPL fitting—such as suboptimal parameter estimation from input images—can propagate errors to subsequent transformations. The model may also hallucinate body parts when target poses differ significantly from input poses. These limitations present opportunities for future work to enhance the fidelity and robustness of human body reshaping in natural images.

### v3

@@ 1 Introduction
Previous methods for human shape editing have relied on 3D morphable models or image warping techniques, which often produce visual artifacts and background distortions under large deformations due to their dependence on explicit geometric correspondences. While human pose editing has advanced significantly with large-scale datasets [20], gathering data of the same person with multiple distinct body shapes presents significant practical challenges. Unlike poses, which can be easily captured in short sessions or extracted from videos, body shape variations are difficult to obtain. This absence of large-scale datasets with identity-consistent body shape variations remains a critical barrier to developing robust shape editing techniques
To address this, we introduce ChangeLing18K, the first large-scale dataset specifically curated for human shape editing. It features diverse body types including thin, muscular, overweight, with each pair depicting the same subject across different shapes while maintaining consistent identity, pose, clothing, and viewpoint. This facilitates precise and identity-preserving model training. Building on this dataset, we propose Odo, to the best of our knowledge, the first ever end-to-end diffusion-based framework designed specifically for human shape editing. Recently, diffusion models have demonstrated remarkable success in different image-editing tasks [23, 7, 22, 16]. Their iterative denoising mechanism, combined with strong priors learned from large datasets, enables fine-grained control and high-quality outputs. However, traditional methods like text-based or point-dragging-based methods [3, 39] are not capable of human body shape transformation as shown in Fig. 2. These techniques lack the precise control for editing human body shapes often causing artifacts or undesired results.
@@
Figure 2: DragDiffusion [39] for fat and thin transformations
@@ 2 Related Work
In this section we discuss the evolution of human-shape editing methods, highlight general diffusion-based techniques, and the current scope of SMPL conditioning in diffusion models.
@@
Image Editing using Diffusion. Foundational work such as Denoising Diffusion Probabilistic Models (DDPM) [12] and Latent Diffusion Models (LDM) [35] introduced scalable frameworks that enable controllable and high-resolution image synthesis. These models have served as the foundation for numerous diffusion-based image-editing works across tasks including inpainting, colorization, semantic modification, and subject personalization.
@@
These methods are designed for semantic, structural, or stylistic editing but are unsuitable for large-scale anatomical transformations in the human body while preserving fine-grained identity and clothing details Fig. 2. Human shape editing requires manipulating body proportions in a controllable manner, beyond the capacity of latent inversion or prompt-conditioned editing.
SMPL Conditioned Diffusion. Recent methods integrate SMPL models into diffusion frameworks to enhance spatial and anatomical consistency in human image generation. CHAMP [48] utilizes SMPL-derived spatial representations as conditioning inputs to a latent diffusion model. By aligning these SMPL conditions with a reference image’s shape parameters, CHAMP achieves temporally coherent human animations. However, its scope is limited to pose variation and does not support body shape editing. Buchheim et al. [4] propose an SMPL-conditioned ControlNet for text-to-image generation, encoding SMPL shape and pose parameters into the diffusion UNet cross-attention. However, their method requires synthetic-to-real domain adaptation and does not facilitate shape editing. These methods show the potential of SMPL conditioning to enhance human shape control within diffusion models.
Thin Fat Muscular
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
3 ChangeLing18K Dataset
@@ 3.1 Dataset Generation Workflow
ChangeLing18K is synthesized through a four–stage, model–driven pipeline that guarantees identity fidelity, pose and attire consistency, and high–resolution realism across three body–types (thin, fat, muscular) Fig. 3. The dataset generation pipeline is shown in Fig. 4 and detailed below.
@@
Figure 5: The left figure illustrates our training pipeline, where the target SMPL depth-map (S) derived from the target image conditions the ReshapeNet via a Depth ControlNet for the desired shape transformation. Features from the reference image, extracted by the ReferenceNet, are integrated into the ReshapeNet using spatial self-attention. The right figure depicts the inference pipeline, where SMPL parameters are initialized from the input image, adjusted based on semantic attributes, and rendered into the target SMPL depth map. This depth map, along with input image features, conditions the ReshapeNet during inference to generate the final output
@@
i) Reference Image selection. The thin image is treated as the baseline reference because it’s smaller silhoutte exposes the most background, yielding a near complete scene for subsequent background completion step.
@@
iv) Uniform Scaling & Composition. From the thin reference mask we measure the height and record the bottom-most coordinates of the mask. Each non-thin mask is resized uniformly to match it’s height, maintaining aspect ratio, and composited onto the clean background at the same coordinates. This ensures consistent framing across images while preserving real shape variations.
@@ 4 Method
Given the above dataset we propose Odo, a diffusion-based approach for reshaping humans. Each pair in the dataset constitutes an input reference image and a target image of a person captured with consistent clothing and background, but a different body shape. Odo incorporates a target SMPL depth map derived from the target image to act as shape conditioning during training. The SMPL model [21] represents the human body’s 3D shape with parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ and pose with joint rotation parameters $\boldsymbol{\theta}\in\mathbb{R}^{72}$. To generate the target depth maps, we first fit a SMPL model to the target image using the method introduced in [37]. The optimized $\boldsymbol{\beta}$ and $\boldsymbol{\theta}$ parameters are subsequently used to render spatially aligned depth maps.
Given an input image $\boldsymbol{I}$ and the target SMPL depth map $\boldsymbol{S}$, our objective is to transform the person in $\boldsymbol{I}$ to match the shape in $\boldsymbol{S}$.
For inference on in-the-wild images without target SMPL depth maps, we use a mapping between semantic attributes and SMPL $\boldsymbol{\beta}$ shape parameters [9]. This mapping enables modification of semantic attributes like weight, height, and body size to control the SMPL $\boldsymbol{\beta}$ shape parameters. Adjusting these parameters modifies the depth map, ensuring the generated output image aligns with the desired body shape.
@@ 5.2 Evaluation
As noted by Okuyama et al. [28], no dataset exists for objective quantitative evaluation of body shape editing with large deformations. We therefore construct the first benchmark dataset for evaluating human body shape editing, consisting of 3,600 image pairs. This benchmark dataset is constructed using the exact same pipeline as the training data but uses distinct real face images and background descriptions to prevent training data leakage. It includes diverse body shape variations for a comprehensive assessment.
We evaluate generated images against the ground truth using four key metrics. For image fidelity, we use Structural Similarity Index Measure (SSIM[41], Peak Signal-to-Noise Ratio (PSNR) for pixel-level accuracy, and Learned Perceptual Image Patch Similarity (LPIPS[46] for perceptual similarity. For body shape accuracy, we fit SMPL models to both ground truth and generated images using [37], then compute the Scale Corrected Per-Vertex Euclidean error in neutral (T-)pose (PVE-T-SC[38]. This metric quantifies how accurately the generated image preserves the target body shape.
@@
In-the-wild Images. We perform an out-of-distribution testing of our model by performing edits on in-the-wild images. Fig. 6 demonstrates this, where our method effectively handles diverse poses, backgrounds, and clothing while preserving person identity. In addition to SMPL target shapes, we provide textual prompts—“Make the person fatter,” “Make the person thinner,” or “Make the person muscular”—to explicitly guide the desired transformations. We have provided more examples in the supplementary material. Fig. 1 further demonstrates our model’s capability to perform diverse shape transformations. The model accurately follows SMPL depth maps to generate multiple variations of thinner and fatter versions from the reference image.
For a quantitative out-of-distribution evaluation we construct a separate dataset using Seedream 4.0 [5]. We collected real world images of people and created fatter, thinnner and muscular versions of them using the Seedream model. Although these edits are not prefect and have some small variations it still serves as a good benchmark to analyse how faithfully Odo is able to follow target SMPL shapes. This dataset consists of 428 image pairs. Our model achieves a low PVE-T-SC of 11.69 mm along with an SSIM of 0.67, PSNR of 18.81, and LPIPS of 0.25. In comparison, the method by Ren et al. achieves an SSIM of 0.62, PSNR of 18.47, LPIPS of 0.26, and PVE-T-SC of 13.82 mm, demonstrating our method’s superior performance.
@@
We also evaluate FLUX.1 Kontext[dev] [2], a prompt-based editing model. We design specific prompts instructing the model to “Make the person fatter,” “Make the person thinner,” or “Make the person muscular” while specifying target weights. However FLUX.1 Kontext[dev] achieves limited performance, as shown in Tab. 2.
Table 2: Comparison of our model Odo against Ren et al. [33] and FLUX.1 Kontext[dev] on our test set. We also show results for three ablations: without prompt conditioning in ReshapeNet, without ReferenceNet, and training only on BR-5K data.
@@ 5.3 Ablations
We conduct ablation experiments to evaluate the effectiveness of various modules in our model. We investigate different methods for inputting image features into ReshapeNet and explore different prompting techniques to guide the model in specific shape transformation tasks. The quantitative results for these ablations are shown in Tab. 2.
@@
We train our model only on the BR-5K dataset and evaluate it on our benchmark data to highlight the significance of our new dataset. Since the BR-5K dataset includes only very slight variations in body shape, the model is unable to learn to perform significant body shape transformations as seen in Fig. 8. This is reflected by the high value of 18.61mm of PVE-T-SC obtained on the test set.
@@ 6 Conclusion
We introduce Odo, the first end-to-end diffusion-based framework for photorealistic, identity-preserving body shape editing at high resolution. Our approach generalizes across diverse images with varying poses and backgrounds by combining a ReferenceNet with a Depth ControlNet-conditioned diffusion model for fine-grained body shape modifications. We also introduce ChangeLing18K, a large-scale dataset containing 18,573 paired images with varied shape transformations while maintaining consistent identity, clothing, and backgrounds.
Figure 9: Failure Case: The target SMPL depth map pose defers significantly from the reference image pose. This may lead to our model hallucinating the original pose in the output
While our approach demonstrates strong performance, there remain areas for improvement. Fine facial details can occasionally be challenging to preserve, leading to facial artifacts. When SMPL target poses differ significantly from input poses, the model may generate inconsistent body parts, as shown in Fig. 9. Fine clothing details (like text) can also be sometimes difficult to reproduce accurately. These limitations present opportunities for future work to enhance fidelity and robustness of human body reshaping.